im trying to get ready for a math test but i dont understand problems like these at all there systems
{y=2x-7
{y=5-x

Found 2 solutions by thechamp1011, indiangangsta:
Answer by thechamp1011(19)

THE SYSTEM OF EUATION MEANS THAT U HAVE TO SOLVE EACH VARIABLE ONE IN TERM OF OTHER...
Y=2X-7....EQ. 1
X=5-Y.....EQ. 2
NOW PLUG IN EQ. 1 IN TO 2
X=5-(2X-7)
X=5-2X+7
X=12-2X
3X=12
X=4....EQ. 3
PLUG IN EQ. 3 INTO EQ. 1
Y=2*3-7=-1
ANS: X=4
AND Y=-1
Y=2(12-2X)-7

Answer by indiangangsta(5)

this problem can use the method of substition. you know that both equations equal to y, so by the transitive property (if a=b, and b=c, then a=c), you can set the two equations equal to each other.
this gives you:
2x-7=5-x
then you solve it as a multistep equation
2x-7=5-x original
3x-7=5 add x to both sides
3x=12 add 7 to both sides
x=4 divide by three
you now know what x is. you must now find out the y point, so that you have a full ordered pair. to do this, just plug in you x value into one of the original equations. Equation two is the easier one, so we will use that
y=5-x original
y=5-4 plug in x
y=1 simplify
thus, your ordered pair is (4,1), and those two lines will meet at that point.
Now to avoid any type of cheating, im gonna type in a bunch of crap, so that, in case you are cheating, you will have to read the whole thing and understand it.
50 cent, young buck, lloyd banks, eminem, and ludacris are good rappers, but the best one will forever be TUPAC SHAKUR. his songs spoke about the hardships of life, and talked about universal themes. in fact, half of his fans were white. not to be racist, but that was what tupac himself said, so im only quoting the great one.
This should be enough, and so i hope u were able to understand and like what i said.